OpenStudy (anonymous):

does (x^2+1) and (x^2-1) consist of imaginary zeros?

OpenStudy (lgbasallote):

hah imaginary zeroes =)) x^2 + 1 seems to have imaginary zeroes though...coz if you equate it as \( x^2 + 1 = 0 \rightarrow x^2 = -1 \rightarrow \sqrt {x^2} = \sqrt {-1}\) do you see where im going?

OpenStudy (anonymous):

yes, i got to sq root of neg one, which is what brings it to an imaginary right.. :) thank youu

OpenStudy (lgbasallote):

right :) and \(\huge \mathsf{\text{ you're welcome}}\)